OpenStudy (anonymous):

LAW OF SINES Find B, C, and c. A = 32°, a = 19, b = 14

OpenStudy (anonymous):

I got B = 23°, C = 125°, c ≈ 29.4. Correct?

OpenStudy (anonymous):

\[\huge \frac{ \sin32 }{ 19 }=\frac{ \sin B }{ 14 } \rightarrow sinB=\frac{ \sin32 }{ 19 } \times 14\]

OpenStudy (anonymous):

\[\huge\ B=\sin^-1\left( \frac{ \sin32 }{ 19}\times14 \right)\]

OpenStudy (anonymous):

\[\huge B= 22.98 \rightarrow B=23°\]

OpenStudy (anonymous):

\[∠C=180-23-32 \rightarrow m∠C=125\]

OpenStudy (anonymous):

\[\frac{ 19 }{ \sin 32 }=\frac{ c }{ \sin125 }\]\[c=\frac{ 19 }{ \sin32 } \times \sin125\]\[\huge c = 29.37°\]

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

You are correct, except that c is a side length so is not measured in degrees. Good job!

OpenStudy (anonymous):

My bad. :D It was a side, not an angle. Thanks!